OpenStudy (anonymous):

Two numbers total 56 and have a difference of 22. Find the twos numbers.

OpenStudy (anonymous):

x+y=56 x-y=22 add and find x

Nnesha (nnesha):

difference means subtract so lets say x and y are two numbers so `Two numbers total 56 ` \[\huge\rm x+y =56\] `difference of 22` \[\huge\rm x-y=22\]now solve for x and y